# GAIA: Rethinking Action Quality Assessment for AI-Generated Videos

### v1

@@ A.1 Ethical Discussions of Our Research
We discuss how our work can be applied to benefit the community. Firstly, the main motivation of our work is that the action-related contents highly affect the video viewing experience, especially in this era where AI-generated models are prevalent, yet current video generation models inevitably suffer from subpar action quality, visual artifacts, and temporal inconsistencies within the generated actions. The existing action quality assessment (AQA) research is highly domain-specific, leading to a relatively poor generalization ability across tasks. Due to the domain gap between real videos and AI-generated videos (AIGVs) as well as the difference in task orientation, previous AQA methods underperform in AI generation scenarios. In terms of data sources, existing AQA studies collect the quality scores directly from judges or minority groups (Tab. 1), which is applicable in professional events but can introduce bias in studying the group preference. The mechanisms by which humans assess the quality of actions and the underlying influences are unknown. In this work, we find that the actions from mainstream T2V models are still subpar in subject quality, action completeness, and action-scene interaction perspectives (even Sora [67] shown in Fig. 7), while neither existing AQA algorithms nor video quality assessment (VQA) methods are suitable for evaluating action quality in AIGVs. Our findings underline the necessity of developing reliable automatic AQA metrics for AIGVs while taking the first step to evaluate the action quality in AIGVs through a causal reasoning manner, which also provides valuable insights for the community in refining video generation models. Secondly, despite the action quality, a common line of works tries to evaluate AI-generated videos from traditional spatial quality (e.g., fidelity, blur, brightness, and aesthetic) and temporal quality (e.g., light change, background consistency, warping error, and motion quality) perspectives [46, 21, 63, 61, 51]. Tab. 6 gives a brief comparison of existing T2V benchmarks. While these lines of work serve a general purpose, their action-related metrics were simply adapted from previous action representation strategies used in real world, which is less effective and exhibits inconsistency with human perception in AI-generated scenarios. Our work helps to build a more reasonable definition of action quality in AIGVs. Thirdly, evaluating action quality in a causal reasoning way offers a promising way to understand human action perception and test the performance of T2V models, thus pointing the path for the future improvement of video generation models.

### v2

@@ 5 Conclusion
6 Acknowledgements
This work was supported in part by the China Postdoctoral Science Foundation under Grant Number 2023TQ0212 and 2023M742298, in part by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20231618, in part by the Shanghai Pujiang Program under Grant 22PJ1407400, and in part by the National Natural Science Foundation of China under Grant 62271312, 62301316, 62101325 and 62101326.
@@ A.1 Ethical Discussions of Our Research
Our work holds the potential for significant social impact, both positive and negative. We anticipate that this work will raise consideration of human perception and understanding in AI-generated actions to better understand generative models and enable more predictable behavior. Currently, the human preference and perceptual sensitivity to the quality of action along the whole action process still remains as an open problem. This work also provides significant guidance on how to optimize video generation models to produce videos with more pleasant actions. Meanwhile, we acknowledge that this study could raise some safety and ethical concerns. One challenging aspect of text-to-video models is the generation of NSFW content (such as violent and pornographic contents), which may be offensive or inappropriate for some viewers and can potentially foster illegal transactions. Although some video generation platforms like MoonValley, Morph and Stable Video have built-in safety filters that detect prompts with NSFW contents, they can still be circumvented through prompt engineering [18]. Additionally, AI video generation technology can be exploited by criminals for fake impersonations and identity theft. Our study also highlighted that some AI-generated videos can convincingly mimic individual’s facial expressions and actions, thereby posing a latent threat to public safety and eroding public trust in social media.
@@
We further discuss how our work can be applied to benefit the community. Firstly, the main motivation of our work is that the action-related contents highly affect the video viewing experience, especially in this era where AI-generated models are prevalent, yet current video generation models inevitably suffer from subpar action quality, visual artifacts, and temporal inconsistencies within the generated actions. The existing action quality assessment (AQA) research is highly domain-specific, leading to a relatively poor generalization ability across tasks. Due to the domain gap between real videos and AI-generated videos (AIGVs) as well as the difference in task orientation, previous AQA methods underperform in AI generation scenarios. In terms of data sources, existing AQA studies collect the quality scores directly from judges or minority groups (Tab. 1), which is applicable in professional events but can introduce bias in studying the group preference. The mechanisms by which humans assess the quality of actions and the underlying influences are unknown. In this work, we find that the actions from mainstream T2V models are still subpar in subject quality, action completeness, and action-scene interaction perspectives (even Sora [72] shown in Fig. 7), while neither existing AQA algorithms nor video quality assessment (VQA) methods are suitable for evaluating action quality in AIGVs. Our findings underline the necessity of developing reliable automatic AQA metrics for AIGVs while taking the first step to evaluate the action quality in AIGVs through a causal reasoning manner, which also provides valuable insights for the community in refining video generation models. Secondly, despite the action quality, a common line of works tries to evaluate AI-generated videos from traditional spatial quality (e.g., fidelity, blur, brightness, and aesthetic) and temporal quality (e.g., light change, background consistency, warping error, and motion quality) perspectives [50, 22, 68, 66, 56]. Tab. 6 gives a brief comparison of existing T2V benchmarks. While these lines of work serve a general purpose, their action-related metrics were simply adapted from previous action representation strategies used in real world, which is less effective and exhibits inconsistency with human perception in AI-generated scenarios. Our work helps to build a more reasonable definition of action quality in AIGVs. Thirdly, evaluating action quality in a causal reasoning way offers a promising way to understand human action perception and test the performance of T2V models, thus pointing the path for the future improvement of video generation models.
@@ B.2 Quantitative and Qualitative Comparison of Content
Figure 14: Scatter plots about MOS against its standard deviation (STD) and five-parameter polynomial fitting plots (orange line) of three perspectives of action quality: (a) subject quality, (b) action completeness, and (c) action-scene interaction.
B.2.1 More Statistics of GAIA
We provide the scatter plots about MOS against standard deviation (STD), along with the five-parameter polynomial fitting plot (orange line) in Fig. 14. First, there is a relatively linear distribution of STD for all three perspectives with MOS<15, suggesting that humans are more consistent in perceiving poor-quality actions. Similar observations can be found in high MOS scenarios (MOS>90). Second, the trend lines reveal a peak in STD distribution when MOS is in $[20,40]$, with a steeper decline and increase in the high MOS range (MOS>80) and low MOS range (MOS<30), respectively. We speculate that AI-generated high-quality actions are mostly consistent with people’s common sense, whereas medium- and low-quality actions exhibit greater diversity, leading to a more pronounced divergence among individuals. Another plausible explanation is that this is due to the uneven distribution of high and low quality action videos in GAIA. Third, the STD distribution is narrower for the subject quality dimension than for action completeness and action-scene interaction dimensions, indicating that the perception of spatial quality distortion in action is less divergent than the temporal consistency and rationality distortion.
@@ C.5 Video-Text Alignment Metrics
Figure 15: Statistical significance comparison among different methods on GAIA dataset. Most p-values are less than 0.001. The methods denoted by ‘1’-‘24’ are USDL, ACTION-NET, CoRe, TSA, Subject Consistency, Motion Smoothness, Dynamic Degree, Human Action, Action-Score, Flow-Score, TLVQM, VIDEVAL, VSFA, BVQA, SimpleVQA, FAST-VQA, DOVER, CLIPScore-ViT-B/16, CLIPScore-ViT-B/32, CLIPScore-ViT-B/32-LAION, CLIPScore-ViT-L/14, BLIPScore, LLaVAScore, and InternLMScore, respectively. Zoom-in for better visualization.
Table 9: 95% confidence intervals (CI) of evaluated methods. Supporting the conclusion obtained in Tab. 5 in the main paper.
Method $\frac{{\overline{SRCC}+\overline{PLCC}}}{2}$ 95% CI
USDL 0.4319 [0.4222, 0.4415]
ACTION-NET 0.4668 [0.4582, 0.4753]
CoRe 0.4456 [0.4373, 0.4538]
TSA 0.4804 [0.4619, 0.4990]
[…]
Motion Smoothness 0.1827 [0.1594, 0.2061]
Dynamic Degree 0.0944 [0.0758, 0.1129]
Human Action 0.2713 [0.2550, 0.2876]
Action-Score 0.2429 [0.2136, 0.2722]
Flow-Score 0.1256 [0.1054, 0.1458]
[…]
VIDEVAL 0.4648 [0.4240, 0.5055]
VSFA 0.5142 [0.4834, 0.5450]
BVQA 0.5186 [0.4835, 0.5537]
SimpleVQA 0.5289 [0.4926, 0.5651]
FAST-VQA 0.5450 [0.5127, 0.5773]
DOVER 0.5534 [0.5161, 0.5908]
[…]
CLIPScoreViT-B/32 0.3735 [0.3540, 0.3930]
CLIPScoreViT-B/32-LAION 0.3402 [0.3259, 0.3545]
CLIPScoreViT-L/14 0.3466 [0.3309, 0.3622]
BLIPScore 0.3913 [0.3654, 0.4172]
LLaVAScore 0.3944 [0.3691, 0.4197]
InternLMScore 0.4124 [0.3883, 0.4365]
Appendix D Extended Results
In this section, we include more observations from the evaluations on the GAIA dataset.
Whether CLIP-based Metrics Excel in Assessing Action Quality? We notice that CLIPScore achieves about 0.38 SRCC and PLCC in the action completeness perspective (Tab. 5), which shows a low correlation with human perception. Although CLIP is not tuned for fine-grained actions, it may work for some coarse-grained actions, as the action itself is also related to the context of scene. We thereby conduct extra experiments to evaluate CLIPScore on three subsets of the GAIA dataset from coarse-grained actions (whole-body) to fine-grained actions (hand and facial). The results are shown in Tab. 10. We can observe that CLIPScore performs significantly worse on the facial subset (an average SRCC of 0.184, 0.194, and 0.239 in terms of subject quality, action completeness, and action-scene interaction, respectively.) than the whole-body (an average SRCC of 0.345, 0.381, and 0.378) and hand subsets. The results further demonstrate the above conjecture that CLIPScore is not appropriate for the assessment of fine-grained actions such as facial actions. Moreover, CLIPScore performs relatively better in action completeness than the subject quality perspective. As discussed in the main paper (Sec. 4.2), we conjecture that such alignment-based metrics are intrinsically sensitive to global high-level vision information (action-related semantics) rather than low-level generative flaws (e.g., blur, noise, textures) that can severely affect the subject quality.
Whether the Combination of Different Metrics can Improve the Perceptual Consistency of Action Quality? We test several different combinations of existing metrics for comparison. As shown in Tab. 11, in most cases, the performance of the combined one is within the best performance of a single one. Surprisingly, we found a performance gain when combining different variants of CLIPScore. We hypothesize that this is due to the spatial feature compensation provided by the different convolutional kernel sizes. Moreover, we observe that combining VSFA with “Human Action” or “Flow-Score” did not yield performance improvements, rather, it resulted in a decrease in SRCC/PLCC scores. We attributed it to different scales of predicted scores, since Flow-Score is an optical flow-based metric. Adding VSFA with three variants of CLIPScore shows better SRCC/PLCC on all three perspectives compared to their single forms. As mentioned in the main paper, VQA methods perform better on subject quality than action completeness and action-scene interaction perspectives, which is opposed to CLIPScore. Therefore, combining these two kinds of metrics could effectively improve the subjective consistency of results. This observation provides intuition for the future development of better AQA methods. Additionally, CLIPScore and its variants outperform the other methods under zero-shot settings. This result suggests that considering both spatial and textual features to better associate visual features with scene descriptions is helpful in predicting action quality.
Indeed, applying a combination of multiple methods is less efficient in practical applications. In the future, we will explore the structure of different models and investigate the possibility of fusing them at the module level in an end-to-end way to better predict the action quality.
Table 10: Performance comparison on coarse-grained actions (whole-body) and fine-grained actions (hand and facial) from GAIA dataset.
Dimension Subset Subject Completeness Interaction
[…]
CLIPScore (ViT-B/16) Whole-body 0.3381 0.3293 0.3732 0.3656 0.3698 0.3557
Hand 0.3167 0.3084 0.3649 0.3564 0.3361 0.3234
Facial 0.2221 0.2326 0.2307 0.2525 0.2711 0.2861
[…]
Hand 0.3835 0.3788 0.4159 0.4139 0.3964 0.3910
Facial 0.1556 0.1596 0.1747 0.1859 0.2175 0.2201
[…]
Hand 0.3392 0.3269 0.3639 0.3499 0.3373 0.3219
Facial 0.1743 0.1806 0.1775 0.1927 0.2294 0.2359
Table 11: Results for the combination of different metrics on the GAIA dataset.
Dimension Subject Completeness Interaction
[…]
Human Action 0.2453 0.2369 0.2895 0.2812 0.2861 0.2743
Action-Score 0.2023 0.1823 0.2867 0.2623 0.2689 0.2432
Flow-Score 0.1471 0.1541 0.0816 0.1273 0.1041 0.1309
Human Action+Action-Score 0.1530 0.1355 0.2333 0.2098 0.2156 0.1912
Human Action+Flow-Score 0.1567 0.1550 0.0940 0.1293 0.1155 0.1324
Action-Score+Flow-Score 0.1199 0.1464 0.0439 0.1175 0.0679 0.1214
Human Action+Action-Score+Flow-Score 0.1279 0.1484 0.0530 0.1198 0.0767 0.1237
[…]
VSFA+Human Action 0.0836 0.0790 0.0059 0.0142 0.0135 0.0096
VSFA+Action-Score 0.2599 0.2531 0.3149 0.3046 0.3054 0.2939
VSFA+Flow-Score 0.1309 0.1506 0.0714 0.1253 0.0914 0.1283
[…]
DOVER 0.6173 0.6301 0.5198 0.5323 0.5164 0.5278
TSA + DOVER 0.5744 0.5831 0.5068 0.5147 0.5081 0.5158
[…]
CLIPScore-B/32 0.3398 0.3330 0.3944 0.3871 0.3875 0.3821
CLIPScore-L/14 0.3211 0.3156 0.3657 0.3574 0.3585 0.3426
CLIPScore-B/16+CLIPScore-B/32 0.3746 0.3698 0.4234 0.4172 0.4148 0.4028
CLIPScore-B/16+CLIPScore-L/14 0.3479 0.3428 0.3967 0.3893 0.3878 0.3738
CLIPScore-B/32+CLIPScore-L/14 0.3747 0.3687 0.4218 0.4145 0.4140 0.3998
CLIPScore-B/16+CLIPScore-B/32+CLIPScore-L/14 0.3734 0.3681 0.4227 0.4157 0.4140 0.4006
[…]
VSFA+CLIPScore-B/32 0.4162 0.4120 0.4377 0.4355 0.4364 0.4288
VSFA+CLIPScore-L/14 0.3651 0.3582 0.3826 0.3793 0.3821 0.3709
VSFA+CLIPScore-B/16+CLIPScore-B/32 0.4004 0.3938 0.4361 0.4303 0.4308 0.4192
[…]